# Geometric Dynamics of Agentic Loops in Large Language Models Trajectories, Attractors and Dynamical Regimes in Semantic Space

### v4

@@ 5.1 Experimental Design
All experiments are performed locally using the Ollama inference framework (version 0.12.9) with the deepseek-r1:8b model. This model provides a lightweight yet expressive transformer architecture, suitable for studying fine-grained textual transformations over multiple recursive steps. We use a sampling temperature of $0.8$, which introduces mild stochasticity while preserving semantic coherence, ensuring that the observed trajectories reflect both the model’s internal tendencies and the intrinsic stability of the prompt formulation. No explicit random seed was specified, so each generation uses one determined by the inference framework, contributing to the stochastic nature of the observed dynamics.
@@
The process is repeated for $T=50$ iterations, generating a trajectory $\{a_{t}\}_{t=0}^{T}$ and its projection $\{e_{t}=\psi(a_{t})\}_{t=0}^{T}$ in the representation space $\mathcal{E}$. All trajectories are embedded using the same function $\psi$ and analyzed in the calibrated similarity space defined in Section 3.
@@
This grid allows both fine-grained and coarse-grained analyses of the clustering structure, ensuring that attractor detection is not an artifact of a single threshold choice.
@@
    The sequence of pairwise displacements $\|e_{t}-e_{t-1}\|_{2}$ and calibrated similarities $\tilde{s}(e_{t},e_{t-1})$, quantifying instantaneous dynamics.
@@
    The evolution of the distance $\|e_{t}-e_{0}\|_{2}$ and similarity $\tilde{s}(e_{t},e_{0})$ to the initial state, capturing long-term semantic drift or contraction.
@@ 5.2.2 Cluster and Attractor Analysis
We now apply the density-based clustering framework defined in Section 4.2 to characterize how the trajectory organizes into stable subregions or attractors. Three clustering configurations were tested, varying only the spatial radius $\rho$ while keeping $\lambda=0.8$ and $\kappa=2$ constant.
Configuration $(\lambda,\rho,\kappa)=(0.8,0.1,2)$.
Cluster Iteration Range Dispersion Outliers
Cluster 1 0–3 0.0648 2
Cluster 2 6–12 0.0400 2
Cluster 3 15–17 0.0400 2
Cluster 4 20–27 0.0967 4
Cluster 5 30–47 0.0800 8
Figure 3 shows the temporal progression of cluster memberships. The transition between clusters follows a smooth temporal sequence rather than abrupt shifts, reflecting progressive refinement.
Attractor similarity matrix.
The inter-cluster similarity matrix (Table 2) reveals strong geometric coherence among all cluster centroids.
C1 C2 C3 C4 C5
C1 1.0000 0.9800 0.8060 0.8580 0.8500
C2 0.9800 1.0000 0.8326 0.7956 0.7804
C3 0.8060 0.8326 1.0000 0.7937 0.8600
C4 0.8580 0.7956 0.7937 1.0000 0.9800
C5 0.8500 0.7804 0.8600 0.9800 1.0000
All off-diagonal similarities remain above $0.78$, consistent with the observation that even when the trajectory transiently visits different micro-regions, these remain tightly aligned in the semantic manifold. Clusters 4 and 5 exhibit the highest cohesion and persistence, suggesting convergence toward a single dominant attractor.
Interpretation.
The contractive loop thus exhibits a robust convergence regime characterized by:
  • bounded local and global geometric variation;
  • smooth temporal progression across increasingly stable clusters;
  • and a high-density attractor region exhibiting inter-cluster similarity above $0.8$.
From a linguistic perspective, this behavior reflects progressive syntactic smoothing and stylistic normalization: the LLM converges toward a canonical phrasing that represents an equilibrium point in the semantic manifold.
Configuration $(\lambda,\rho,\kappa)=(0.8,0.2,2)$.
When the clustering radius is increased to $\rho=0.2$, smaller micro-clusters merge into broader regions, revealing a simplified attractor landscape composed of only two persistent clusters. These clusters correspond to two major stabilization phases within the contractive trajectory: an initial transient phase and a final convergence basin.
@@
The temporal organization of the trajectory, shown in Figure 4, confirms a single transition between two stable basins. The early iterations (Cluster 1) exhibit slightly higher dispersion, while the second phase (Cluster 2) displays long persistence and near-perfect internal similarity.
@@
Attractor similarity matrix.
The inter-cluster similarity matrix in Table 4 shows that both clusters are strongly aligned in the embedding space, with a centroid similarity of $0.89$.
C1 C2
C1 1.0000 0.8900
C2 0.8900 1.0000
The high centroid similarity confirms that the two attractors occupy closely neighboring regions in $\mathcal{E}$, consistent with a smooth convergence process. The transition between them occurs around iteration 20 and marks the point where the rewriting process reaches its asymptotic stability.
Interpretation.
Compared to the previous configuration ($\rho=0.1$), this setting highlights the macro-scale structure of the trajectory: the loop first explores a transient manifold (Cluster 1) before collapsing into a single dense attractor (Cluster 2). The overall trajectory thus reflects a monotonic contraction process toward a stable equilibrium region, with no evidence of oscillation or divergence.
Configuration $(\lambda,\rho,\kappa)=(0.8,0.3,2)$.
At a larger radius $\rho=0.3$, the clustering process identifies a single dense region encompassing the entire trajectory. This configuration effectively merges all transient substructures into one unified attractor, consistent with the strongly contractive nature of the rewriting loop.
Cluster Iteration Range Dispersion Outliers
Cluster 1 0–50 0.2100 13
The corresponding temporal distribution, shown in Figure 5, confirms the absence of any segmentation: all iterations are assigned to the same cluster, with a few boundary outliers reflecting minimal fluctuations near the threshold $\rho=0.3$.
Interpretation.
The detection of a single persistent cluster across all iterations confirms that the rewriting operator $F$ acts as a strong contraction mapping in (,s~)($\mathcal{E}$$\tilde${s}). Even under relaxed density constraints, no internal bifurcation or oscillatory pattern emerges. This configuration thus captures the fixed-point regime of the loop, where the generated sentences become semantically and stylistically indistinguishable over time.
Summary of clustering outcomes.
When the clustering radius $\rho$ increases to $0.2$ or above, micro-clusters merge into a single dense region, effectively collapsing the entire trajectory into one stable attractor. This collapse into a single dense attractor empirically validates the $\mathcal{R}_{\mathrm{ctr}}$ definition (Section 4.5), suggesting convergence toward a stable fixed point in the embedding space. No secondary structure is detected for $\rho\geq 0.2$, consistent with a globally contractive process with negligible oscillation. However, these observations are based on $T=50$ iterations. While the trajectory shows no signs of leaving the attractor within this window, we cannot formally guarantee that oscillations or cluster exits would not occur over longer horizons. The empirical stability observed here suggests such escapes are unlikely, but proving this rigorously would require a more advanced mathematical framework, for instance analyzing the probability of entering and remaining within a cluster as a function of the transformation operator $F$ and the embedding geometry.
@@ 5.3 Exploratory Loop Dynamics
In contrast to the previous rewriting loop, the exploratory loop was designed to alternate between summarization and semantic negation. This alternation introduces an intrinsic instability in the semantic space, continuously disrupting coherence and preventing any convergence toward a stable region of $\mathcal{E}$. The resulting trajectory exhibits high variability and recurrent semantic inversions, typical of a divergent or oscillatory dynamic. The complete sequence of generated artifacts is provided in Appendix B for reference.
@@ 5.4 Comparative Summary
These results provide empirical evidence for the existence of distinct dynamic regimes in recursive generative systems. The contractive loop exhibits progressive semantic refinement, converging toward a stable attractor as evidenced by high local similarity (>0.85>0.85) and bounded global drift. In contrast, the exploratory loop exhibits persistent instability, with each iteration producing semantically distant outputs ($\sim$1.0 Euclidean distance) and no detectable cluster structure. This dichotomy confirms that the choice of prompt and transformation operator fundamentally determines the long-term behavior of agentic trajectories, ranging from stable convergence to unbounded semantic wandering.
@@ 6.3 Toward Geometry of Agentic Cognition
The trajectory analysis framework developed in this paper enables a natural question: can an agentic loop reason about its own geometric dynamics and adapt its behavior to achieve specific objectives? This motivates a future research direction we term Geometry of Agentic Cognition: systems where the agent possesses geometric self-awareness and can metacognitively control its trajectory evolution.
Current state: external control.
Presently, trajectory control remains external to the loop itself. The intensity of contraction or exploration can be modulated through:
  • Prompt engineering: adjusting the degree of rewriting, abstraction or negation in the instruction layer;
  • Sampling parameters: manipulating temperature, nucleus probability (top-p), top-k or repetition penalties to alter stochasticity.
These mechanisms require external intervention: a human engineer or outer control system must specify parameters a priori.
Future direction: agentic geometric cognition.
A fundamentally different paradigm would enable the agentic loop itself to observe, reason about and adapt its trajectory dynamics based on geometric awareness. In this vision, at each iteration $t$, the agent would:
  1. 1.
    Observe trajectory geometry: Compute calibrated similarities $\tilde{s}(e_{t},e_{t-1})$ (local dynamics), $\tilde{s}(e_{t},e_{0})$ (global drift) and trajectory dispersion to assess its current dynamical regime.
  2. 2.
    Reason about objective alignment: Evaluate whether its trajectory is progressing toward the desired objective, whether convergence toward a solution (requires contraction), exploration of alternatives (requires divergence) or escape from a suboptimal attractor (requires controlled instability).
  3. 3.
    Adapt iteration strategy: Based on this geometric self-assessment, metacognitively adjust the next transformation by increasing exploration if trapped in local minima, inducing contraction if approaching the target region or maintaining the current regime if progressing optimally.
This would transform agentic loops from passive trajectory executors into geometric navigators: systems that understand their position in semantic space and actively steer toward objectives through informed dynamical control.
Enabling mechanisms.
Implementing agentic geometric cognition requires:
  • Trajectory introspection: Embedding the agent’s own outputs $\{a_{0},a_{1},\ldots,a_{t}\}$ and computing geometric indicators (similarity, dispersion, clustering metrics) as part of the loop context.
  • Geometric prompt characterization: Pre-characterizing prompts by their dynamical properties (contractive, exploratory) and selecting from a prompt library based on desired trajectory behavior. This requires a mathematical framework capable of predicting a prompt’s dynamical regime a priori, which we address in forthcoming work.
  • Adaptive parameter control: Allowing the agent to specify its own generation parameters or select from predefined strategies (e.g., "use high-temperature exploration" vs. "use low-temperature refinement") based on trajectory analysis.
This research direction opens fundamental questions at the intersection of dynamical systems, metacognition and autonomous reasoning: Can language models learn to recognize geometric patterns in their own trajectories? Can they develop heuristics for when to contract versus explore based on objective alignment? Can they generalize trajectory control strategies across different tasks? The geometric framework established in this paper provides the measurement infrastructure required to empirically investigate these questions, enabling rigorous quantification of whether and how agentic systems can achieve geometric self-regulation.
@@ Appendix A Contractive Loop: Complete Trajectory
Appendix B Exploratory Loop: Complete Trajectory

### v5

@@ 1 Introduction
Claims investigated.
  • C1: Singular loops induced by prompt P can fall into regimes measurable in the embedding space.
  • C2: Under a paraphrase prompt, trajectories form persistent clusters with identifiable attractors and decreasing dispersion over time.
  • C3: Under a negation prompt, no persistent cluster emerges under the tested similarity and dispersion thresholds.
@@ 4.5 Dynamic Regimes
4.6 Assumptions and Scope
The framework introduced in this work relies on a number of explicit assumptions and scope restrictions, which we state here to clarify the validity and intended interpretation of the proposed definitions and empirical observations.
A1. Black-box transformation operator.
We assume that the transformation operator $F:\mathcal{A}\rightarrow\mathcal{A}$, implemented via a prompt and a language model invocation, is treated as a black box. No assumptions are made about its internal architecture, differentiability, Lipschitz continuity, or contractive properties in the classical analytical sense. All characterizations of dynamics are therefore observational rather than analytical.
A2. Finite-horizon trajectories.
Agentic trajectories are observed over a finite number of iterations $T$. All notions of convergence, divergence, and stability are defined operationally with respect to finite trajectories, rather than as asymptotic guarantees. The framework does not claim formal convergence in the limit $T\rightarrow\infty$.
A3. Representation-based observation.
While agentic loops operate in the discrete artifact space $\mathcal{A}$, their dynamics are observed indirectly through a projection into a continuous representation space $\mathcal{E}$ via an embedding function $\psi$. We assume that relative semantic proximity in $\mathcal{E}$ reflects meaningful semantic relationships between artifacts, but make no claim that $\mathcal{E}$ constitutes the true state space of the generative process.
A4. Order-preserving similarity measures.
The proposed definitions of clusters, attractors, and dynamical regimes rely on relative similarity comparisons rather than absolute metric values. Consequently, the framework assumes only that the similarity function preserves the ordering of semantic proximity. While calibrated similarity is used for interpretability, the qualitative classification of regimes is invariant under any monotonic transformation of the similarity measure.
A5. Prompt-defined operators and stochasticity.
The effective transformation operator $F$ is defined jointly by the prompt template and the language model decoding configuration. Sampling stochasticity is therefore an inherent component of the dynamics. Observed regimes characterize the statistical behavior induced by a given prompt-model configuration, rather than deterministic properties of the model alone.
These assumptions delineate the scope of the framework. The objective is not to derive formal guarantees about language model dynamics, but to provide an operational and interpretable methodology for characterizing and comparing the observable behavior of iterative agentic loops.
@@ 5.1 Experimental Design
All experiments are performed locally using the Ollama inference framework (version 0.12.9) with the deepseek-r1:8b model. This model was selected for three reasons: (i) it is small enough to run locally without GPU clusters, enabling full reproducibility; (ii) despite its size, it demonstrates strong instruction-following capabilities on rewriting and reasoning tasks, ensuring that prompt-induced behaviors are faithfully executed; and (iii) local execution eliminates API rate limits and guarantees deterministic infrastructure, which is essential when running 50-iteration loops repeatedly. We use a sampling temperature of $0.8$, which introduces mild stochasticity while preserving semantic coherence, ensuring that the observed trajectories reflect both the model’s internal tendencies and the intrinsic stability of the prompt formulation. No explicit random seed was specified, so each generation uses one determined by the inference framework, contributing to the stochastic nature of the observed dynamics.
@@
The process is repeated for $T=50$ iterations, generating a trajectory $\{a_{t}\}_{t=0}^{T}$ and its projection $\{e_{t}=\psi(a_{t})\}_{t=0}^{T}$ in the representation space $\mathcal{E}$. This horizon balances computational cost against sufficient observation length: preliminary runs indicated that contractive loops stabilize within 20–30 iterations, while exploratory loops show no sign of convergence; $T=50$ thus provides adequate margin to distinguish these regimes. All trajectories are embedded using the same function $\psi$ and analyzed in the calibrated similarity space defined in Section 3.
@@
The similarity threshold $\lambda=0.8$ enforces high local coherence, admitting only consecutive pairs that are clearly semantically related. The dispersion radii $\rho\in\{0.1,0.2,0.3\}$ provide multi-resolution analysis: $\rho=0.1$ detects only tightly packed micro-clusters, while $\rho=0.3$ captures broader attractor basins; the increments are chosen to span interpretable fractions of the calibrated similarity scale. The patience parameter $\kappa=2$ tolerates brief noise (up to two consecutive violations) without prematurely terminating a cluster, while remaining strict enough to separate genuine regime transitions. This grid allows both fine-grained and coarse-grained analyses of the clustering structure, ensuring that attractor detection is not an artifact of a single threshold choice.
@@
    The sequence of calibrated similarities $\tilde{s}(e_{t},e_{t-1})$, quantifying instantaneous dynamics.
@@
    The evolution of the similarity $\tilde{s}(e_{t},e_{0})$ to the initial state, capturing long-term semantic drift or contraction.
@@ 5.2.2 Cluster and Attractor Analysis
We now apply the density-based clustering framework defined in Section 4.2 to characterize how the trajectory organizes into stable subregions or attractors. Three clustering configurations were tested, varying only the spatial radius $\rho\in\{0.1,0.2,0.3\}$ while keeping $\lambda=0.8$ and $\kappa=2$ constant. We present the analysis for $\rho=0.2$, which provides an intermediate resolution that clearly reveals the attractor structure; complete results for all configurations are provided in Appendix B.
Cluster structure at $\rho=0.2$.
At this resolution, the clustering algorithm identifies two persistent clusters corresponding to distinct stabilization phases: an initial transient phase (iterations 0–18) and a final convergence basin (iterations 21–50).
@@
The temporal organization of the trajectory, shown in Figure 3, confirms a single transition between two stable basins around iteration 20. The early iterations (Cluster 1) exhibit slightly higher dispersion reflecting initial semantic exploration, while the second phase (Cluster 2) displays long persistence and decreasing internal dispersion—hallmarks of convergence toward a stable attractor.
@@
Multi-resolution robustness.
Identical qualitative behavior is observed across all tested resolutions: at $\rho=0.1$ (finer resolution), five micro-clusters emerge that merge smoothly over time; at $\rho=0.3$ (coarser resolution), a single cluster encompasses the entire trajectory. In all cases, clusters exhibit temporal persistence, high inter-cluster similarity (>0.78>0.78), and monotonically decreasing dispersion. See Appendix B for complete tables, figures, and similarity matrices.
Summary.
The cluster analysis confirms that the contractive loop satisfies the empirical criteria for $\mathcal{R}_{\mathrm{ctr}}$ (Section 4.5):
  • persistent clusters with identifiable attractors;
  • decreasing dispersion indicating convergence;
  • no oscillation or divergence across 50 iterations.
The consistency of these findings across three clustering resolutions—without parameter tuning—strengthens the conclusion that the observed convergence reflects genuine geometric properties of the trajectory rather than artifacts of a particular configuration. However, these observations are based on $T=50$ iterations; while the trajectory shows no signs of leaving the attractor within this window, proving long-term stability rigorously would require analyzing the probability of entering and remaining within a cluster as a function of the transformation operator $F$ and the embedding geometry.
@@ 5.3 Exploratory Loop Dynamics
In contrast to the previous rewriting loop, the exploratory loop was designed to alternate between summarization and semantic negation. This alternation introduces an intrinsic instability in the semantic space, continuously disrupting coherence and preventing any convergence toward a stable region of $\mathcal{E}$. The resulting trajectory exhibits high variability and recurrent semantic inversions, typical of a divergent or oscillatory dynamic. The complete sequence of generated artifacts is provided in Appendix C for reference.
@@ 5.4 Comparative Summary
These results provide empirical evidence for the existence of distinct dynamic regimes in recursive generative systems. The contractive loop exhibits progressive semantic refinement, converging toward a stable attractor as evidenced by high local similarity (>0.85>0.85) and bounded global drift. In contrast, the exploratory loop exhibits persistent instability, with each iteration producing semantically distant outputs ($\sim$1.0 Euclidean distance) and no detectable cluster structure. These findings substantiate Claims C1–C3. This dichotomy confirms that the choice of prompt and transformation operator fundamentally determines the long-term behavior of agentic trajectories, ranging from stable convergence to unbounded semantic wandering.
@@ 6 Discussion and Perspectives
6.1 Limitations
The present study has several limitations that scope its conclusions:
  • All experiments use a single LLM (deepseek-r1:8b) with fixed decoding settings (temperature $0.8$); different architectures or sampling strategies may yield different dynamical signatures.
  • Trajectories are observed through a single embedding function $\psi$ and calibrated similarity $\tilde{s}$; alternative embeddings or metrics could reveal different geometric structures.
  • The observation horizon is limited to $T=50$ iterations, which may not capture slower convergence phenomena or late-emerging instabilities.
  • The two prompts were chosen as archetypal extremes (pure paraphrase, pure negation) to cleanly separate regimes; they do not constitute an exhaustive map of the space of possible dynamical behaviors.
@@ 6.4 Toward Geometry of Agentic Cognition
The trajectory analysis framework developed in this paper enables a natural question: can an agentic loop reason about its own geometric dynamics and adapt its behavior accordingly? Currently, trajectory control remains external—modulated through prompt engineering or sampling parameters (temperature, top-p) specified a priori by a human engineer.
A fundamentally different paradigm, which we term Geometry of Agentic Cognition, would enable the loop itself to observe its trajectory geometry (local similarity, global drift, dispersion), reason about whether its current regime aligns with the objective (convergence toward a solution vs. exploration of alternatives), and adapt its strategy accordingly. This would require trajectory introspection (embedding recent outputs and computing geometric indicators within the loop context) and geometric prompt characterization (pre-classifying prompts by their dynamical properties to enable informed selection). Such systems would transform from passive trajectory executors into geometric navigators that actively steer through semantic space. The framework established here provides the measurement infrastructure for investigating whether agentic systems can achieve such geometric self-regulation.
@@ Appendix A Contractive Loop: Complete Trajectory
Appendix B Robustness of Contractive Dynamics to Clustering Parameters
This appendix provides complete clustering results for the contractive loop across all tested configurations, demonstrating that the qualitative conclusions of Section 5.2 are robust to the choice of clustering resolution $\rho$.
Configuration $(\lambda,\rho,\kappa)=(0.8,0.1,2)$
Cluster Iteration Range Dispersion Outliers
Cluster 1 0–3 0.0648 2
Cluster 2 6–12 0.0400 2
Cluster 3 15–17 0.0400 2
Cluster 4 20–27 0.0967 4
Cluster 5 30–47 0.0800 8
The inter-cluster similarity matrix (Table 4) reveals strong geometric coherence among all cluster centroids, with all off-diagonal similarities above $0.78$.
C1 C2 C3 C4 C5
C1 1.0000 0.9800 0.8060 0.8580 0.8500
C2 0.9800 1.0000 0.8326 0.7956 0.7804
C3 0.8060 0.8326 1.0000 0.7937 0.8600
C4 0.8580 0.7956 0.7937 1.0000 0.9800
C5 0.8500 0.7804 0.8600 0.9800 1.0000
Configuration $(\lambda,\rho,\kappa)=(0.8,0.2,2)$
The inter-cluster similarity matrix for the intermediate resolution (main text, Table 1) shows that both clusters are strongly aligned in the embedding space, with a centroid similarity of $0.89$.
C1 C2
C1 1.0000 0.8900
C2 0.8900 1.0000
Configuration $(\lambda,\rho,\kappa)=(0.8,0.3,2)$
At a larger radius $\rho=0.3$, the clustering process identifies a single dense region encompassing the entire trajectory, effectively merging all transient substructures into one unified attractor.
Cluster Iteration Range Dispersion Outliers
Cluster 1 0–50 0.2100 13
Summary Across Configurations
Table 7 consolidates the clustering outcomes across all three resolutions.
Resolution $\rho$ Clusters Min Similarity Interpretation
0.1 (fine) 5 0.78 Micro-phases of refinement
0.2 (medium) 2 0.89 Transient + convergence basin
0.3 (coarse) 1 Single unified attractor
Table 7: Summary of clustering outcomes for the contractive loop across three resolution levels.
All configurations reveal the same qualitative pattern: temporally persistent clusters with high inter-cluster similarity and no oscillation or divergence. The number of detected clusters decreases monotonically with increasing $\rho$, as expected from hierarchical structure in the embedding space. This consistency across resolutions confirms that the observed convergence is a genuine property of the trajectory rather than an artifact of parameter choice.
Appendix C Exploratory Loop: Complete Trajectory